# Robust shallow shadows

### v1

@@ Results.
We begin by assessing the need of mitigating noise in shallow circuits. The top panel of Fig. 2 shows the worst-case bias, $\max_{O,\varrho}\,\operatorname{bias}(O,\varrho)$, of standard non-mitigated shadow estimation for Pauli observables (see App. B for details) for different system sizes, circuit depths, and incoherent noise levels. The worst-case bias is monotonically increasing with all three parameters and the results are consistent with a linear increase of the bias with both the circuit depth and system size. For two-layer shadows, we observe Pauli worst-case biases that are already for three to four qubits about one order of magnitude larger than the average infidelity of the CNOTs used in the circuit. For example, targeting an estimation precision of $10^{-2}$, even fairly accurate CNOTs gates with infidelity $10^{-3}$ can introduce a sizeable biases in the linear estimation. The stability is improved for Pauli observables with bounded support, see App. D. Nonetheless, without mitigation using even very shallow shadows is often not possible already for small system sizes. This underpins the motivation to study our noise-robust variant.
@@
A key application of shadows with global Clifford rotations is fidelity estimation. Here, the sample complexity is constant in the system size, while randomized local bases measurements have exponential scaling [16]. This leads to the question whether a very shallow hardware-efficient ansatz can improve the sample efficiency even in the presence of experimental imperfections. We simulate the protocol of robust shadow fidelity estimation assuming perfectly prepared pure input states. The fidelity is estimated with respect to the input state itself. Only the two-qubit gate implementing the basis rotation for the shadow suffer from incoherent noise with average infidelity $10^{-3}$ and the number of samples is fixed to |$\absolutevalue{\mathcal{S}}=10^{5}$. The noisy frame operator is estimated with |$\absolutevalue{\mathcal{S}}=10^{6}$, which we found sufficient for ensuring that the estimation error of $\tilde{S}$ does not limit the accuracy of the fidelity estimation.
@@
Fig. 2 (bottom) shows the estimation error averaged over multiple uniformly-random pure states for different system sizes and circuit depths. We find that the system-size dependence is significantly reduced already for shadows with $D=1,2$ when compared to local circuits. In effect, the estimation error is reduced by a factor of about $2$ already for these small systems of size $n=7,8$ when using short-depth circuits. Importantly, we observe that, in order to harness these effects, the empirical estimation of $\tilde{S}$ put forward here is essential. Without noise-mitigation, the noise-induced bias, depicted in the figure with dashed lines, deteriorates fidelity estimation by factors of $3$ to $4$ for the same small system sizes and $D=1$. For $D=2$ the estimation error using the noiseless frame operator $S$ is already almost one order of magnitude worse than the one that uses $\tilde{S}$.
@@
In general, even assuming perfect bias mitigation, improving the estimation error by going to deeper circuits will depend on the infidelity of the CNOT gates. Intuitively, every CNOT with incoherent errors will increase the probability that one only learns compromised information about the state. Even if the noise is perfectly mitigated by accurately estimating $\tilde{S}$, with some probability we learn less about the state. Thus, there will be an increased sample complexity of the fidelity estimation itself. At the same time, longer circuits (without noise) will improve the variance of fidelity estimation eventually reaching the efficacy of $3$-designs. We therefore have two competing effects on the sample complexity. We expect that for many estimation tasks there exists a cross-over point depending on the infidelity of the entangling gates. Below this cross-over infidelity, increasing the depth improves estimation accuracy (at fixed number of samples). Above it, accuracy is deteriorated by deeper circuits.
@@
Figure 3: Tensor network representation of robust measurement frames of shallow circuits. (a) Worst-case bias for the estimation of Pauli observables introduced by converting $\tilde{S}$ to a matrix product state (TT) representation using sequential SVDs, relative to the size of the corresponding eigenvalue, as a function of $n$, for different bond dimensions $\chi$, and depths $D=1,2$. We observe that TT ranks $4$ and $8$ suffice to describe $\tilde{S}$ up-to statistical error. (b) Worst-case bias for the estimation of Pauli observables when learning $\tilde{S}$ directly from the calibration data using a modified ALS method. We find a similar performance as for the sequential SVD.
We now take a more detailed look at the TT structure of the frame operator and its inverse. We begin with using sequential SVDs to solve the recovery problem in Eq. (6) for different ranks $\chi$ using the same noise setting as above. Fig. 3 panel (a) shows the additional (relative) worst-case bias in estimating Pauli observables when using a TT approximation to $f$ instead of the empirically learned dense vector. We indeed find that for $D=1$ a rank $\chi=4$ approximation yields a small relative error that is constant with the system size for $n\geq 6$ and higher ranks do not yield an improvement. Furthermore, we find that for smaller system-sizes the MPS is also fitting the statistical fluctuations. For $D=2$ we find that $\chi=8$ accurately approximates $\tilde{S}$ up to statistical error.
We then apply the computationally and storage-efficient MALS algorithm to directly learn a TT representation of $\tilde{S}$ from the shadow data and calculate the induced bias in estimating Pauli observables. We initialize the algorithm with the analytical result for $S$ for the local Clifford frame. We observe a identical performance to the SVD. MALS is rank-adaptive but does not increase the rank beyond $\chi=8$ for $D=2$ even if permitted.
Figure 4: Simulation of overlap estimation from Ref. [5] under a gate-dependent noise model. (left) Worst-case bias for the estimation of Pauli observables introduced by implementing short-depth circuits with and without noise in an experiment, but using the ideal global Clifford frame in the post-processing stage, as a function of $D$, for two frames: $S_{n=4}^{\otimes 2}$ (noiseless, blue line) and $\tilde{S}_{n=4}^{\otimes 2}$ (noisy, brown line). Frames were estimated with |$\absolutevalue{\mathcal{S}}=10^{6}$. Noisy frame was generated under a gate-dependent circuit noise as well as readout noise. (right) Fidelity estimation error using aforementioned frames. States |$\ket{\Psi_{T}}$, |$\ket{\phi_{1}}$, and |$\ket{\phi_{2}}$ are defined in Appendix E.
@@
Quantum hardware demonstration.
In Fig. 5, we show the results of an experimental implementation of robust shallow shadows on the ibm_nairobi quantum processor from IBMQ. We implemented the measurement circuit of Fig. 1 in $5$ of the $7$ qubits available in the device, for depths $D\in\{0,1,2\}$. In Fig. 5 (a) we plot the estimated eigenvalues of $\tilde{S}$ as a function of the cardinality of the shadow $|\mathcal{S}|$ up to |$\absolutevalue{\mathcal{S}}=10^{5}$. (We also provide histograms for the distribution of samples in Fig. 4 (a) in App. F.) For simplicity, we show one eigenvalue per Pauli support. The dashed lines show the eigenvalues of the corresponding noiseless frame. In all three cases, the eigenvalues differ from their noiseless counterparts by a margin that is statistically significant, This confirms that the measurement frame to be inverted in practice should indeed be $\tilde{S}$ instead of $S$. To cross-validate our results, we separate our $10^{5}$ samples in two batches: with a first batch of $90\%$ of the snapshots, we estimate $\tilde{S}$; with the remaining $10\%$, in turn, we perform a shadow tomography protocol on the 5-qubit state |$\outerproduct{\mathbf{0}}{\mathbf{0}}$ | as well as on its marginal state |$\outerproduct{0}{0}$ | on one the qubits. For each of the two states, we invert the classical shadows using both $\tilde{S}^{-1}$ and $S^{-1}$. In Fig. 5 (b), we plot the fidelity biases introduced by the mitigated and unmitigated estimations, respectively. For the unmitigated estimation, one can observe that at |$\absolutevalue{\mathcal{S}}\approx 10^{3}$ the statistical error becomes smaller than the bias (i.e., the systematic error due to the ‘wrong’ inverted frame). Increasing the number of samples does not improve the accuracy. In contrast, for the robust shallow shadows, the accuracy continues to increase steadily as |$\absolutevalue{\mathcal{S}}$ | goes beyond $10^{3}$.
Conclusions.
We presented a robust classical shadow scheme for a broad family of noisy, shallow measurement circuits. The measurement calibration stage admits an efficient matrix-product formulation, and the robust inversion stage can also be approximated using standard tensor-network tools. We illustrated the relevance of our method with both numerics and proof-of-principle experiments on an IBM Q device. Numerically, for relevant parameter regimes, we showed that while unmitigated shallow shadows with noisy circuits become more biased as the depth increases, robust ones become significantly more accurate up-to a crossing point. Experimentally, in turn, we observed major bias reductions in fidelity estimations using the mitigated variant. Our framework does not require analytical calculations of the effective measurement map but replaces it with a direct estimation protocol. This offers versatile, practical estimation protocols that realize the efficacy of classical shadow as much as it is possible on actually existing and near-term quantum devices by using the available hardware-efficient noisy shallow circuits.
Note added.
During the completion of this paper, another independent work [37] appeared that also studies robust shallow shadows.
Acknowledgements.
We thank Stefano Carrazza and Stavros Efthymiou for their insights on numerical simulations. We further thank Michał Oszmaniec for discussions on noise mitgation. We thank Jadwiga Wilkens for the availability of the Quantum Circuit Library [38].
Appendix A Incoherent noise model
@@
Appendix B Worst-case bias measure
@@
where we used diagonality of $\tilde{S}$ in Pauli basis and that $|(w_{a}\,|\,\varrho)|\leq\|w_{a}\|_{\infty}\|\varrho\|_{1}\leq 1/\sqrt{d}$ for all $a,\varrho$, and $\|\cdot\|_{1}$ denoting the trace-norm (nuclear / Schatten-$1$ norm). Most importantly, this bound is saturated for $\varrho$ in an eigenstate of $w_{a}$ the observable suffering the maximal bias.
Appendix C Details on tensor network description
$\begin{split}\operatorname{CNOT}&=\sum_{i\in\{0,1\}}\outerproduct{i}{i}\otimes X^{i}\\ &=\left(\sum_{i\in\{0,1\}}\outerproduct{i}{i}\otimes\bra{i}\otimes\mathbbm{1}\right)\left(\mathbbm{1}\otimes\sum_{j\in\{0,1\}}\ket{j}\otimes X^{j}\right)\\ &=(L\otimes\mathbbm{1})(\mathbbm{1}\otimes R)\,,\end{split}$ (9)
where $X$ denotes the Pauli-$X$ gate, and the last line defines the operators $L$ and $R$. Now we define the MPO core $C_{i,j}\coloneqq(\bra{i}\otimes\mathbbm{1})R\,L(\mathbbm{1}\otimes\ket{j})\in\mathbb{C}^{2\times 2}$. The unitary operator $L_{\operatorname{CNOT}}$ of a $1$D layer of CNOT gates on $n$ qubits with periodic boundary conditions and even $n$ can be written as
[…]
The operator $\tilde{S}$ can also be written as a shallw circuit of operators acting on quantum channels. Assuming local invariance, also $f=(f_{k})_{k\in[n]}\in\mathbb{R}^{n}$ has a corresponding circuit representation. The rank of the resulting TT description of $f$ is increased by every layer of entangling gates and non-local noise. For fairly general noise affecting the shallow circuits, $f$ is of low TT rank. To this end, it is instructive to discuss a scenario with strictly local and mostly gate-dependent noise. Let us assume that $\Omega^{(i)}_{j}(U)$ is the CPT map describing the noisy implementation of the gate $U\in\mathrm{U}(2)$ when applied on qubit $i$ in the circuit layer $j$. This noise model is general enough to also capture local noise and errors affecting the read-out and entangling gates. These noise effects can be included in the respective neighbouring layer of single-qubit rotations. The local channel twirl with a group $G$ including noise is formally described by the non-commutative Fourier transform of the map $\Omega^{(i)}_{j}$ evaluated at the representation of unitary channels $\omega(U)=U\otimes\bar{U}$, i.e.
$\hat{\Omega}^{(i)}_{j}[\omega]=\int_{G}\bar{U}\otimes U\otimes\Omega^{(i)}_{j}(U)\mathrm{d}\mu(U)\,.$ (11)
We will here simply use $\hat{\Omega}^{(i)}_{j}[\omega]$ as convenient symbol for the “noisy channel twirl”. Note that $\hat{\Omega}^{(i)}_{j}[\omega]$ is a linear operator on superoperators, a “super-duper-operator” if one insists. We can also introduce a corresponding operator for the entangling layer that conjugates a quantum channel with unitary channels of layer of CNOT gates. The resulting MPO, $\mathcal{C}=\bar{L}_{\operatorname{CNOT}}\otimes L_{\operatorname{CNOT}}\otimes L_{\operatorname{CNOT}}\otimes\bar{L}_{\operatorname{CNOT}}$, has bond dimension $16$. The read-out is locally described by $\mathcal{B}_{2}=\sum_{z\in[2]}|\Pi_{z})(\Pi_{z}|$, the local dephasing channel in computational basis. Finally, to ensure the local invariance of $\tilde{S}$, we assume gate-independent left noise $\Omega^{(1)}_{j}(U)=\Lambda_{j}\omega(U)$ for the first layer acting on the state under consideration, with $\Lambda_{j}$ being a quantum channel.
Figure 7 depicts the resulting expression for $f$. Without any further assumption, we conclude that $f$ has TT rank of at most $16^{D}$. The rank is even smaller if all noise channels are Pauli noise. Then, we can fully restrict the circuit in Fig. 7 as acting only on the subspace of Pauli noise channels, i.e. diagonal channels in Pauli operator basis. To see this note that the local dephasing channel is a Pauli noise channel. Under the assumption of Pauli noise, the noisy channel twirl maps Pauli noise channels to Pauli noise channels. Being a Clifford unitary, the entangling layer acts as a two-local permutation on the diagonal of the Pauli-noise channel. The representation of $\mathcal{C}$ restricted to Pauli-noise channels has (MPO) bond dimension $4$. Hence, $f$ is described by an TT of rank at most $\chi=4^{D}$.
The argument for one-local noise can be generalized to quasi-local noise by introducing, e.g., MPO descriptions of the implementation map modelling cross-talk effects, or modelling a noisy entangling layer $\mathcal{C}$ with a potentially larger bond dimension. In general, the TT rank of $f$ will be increased due to correlated noise effects. If these correlations are sufficiently short-ranged, $f$ still has an effective low-rank description. We expect that this will often be the case on actual quantum hardware.
Figure 7: Tensor network of $f$ for strictly local noise mostly gate-dependent noise. The entangling layer acting here as a ‘super-duper operator’ on quantum channels has MPO bond-dimension $16$, restricted to Pauli noise channels it has MPO bond dimension $4$. Correspondingly, the TT rank of $f$ is $16^{D}$ for local noise and $4^{D}$ for local Pauli noise.

### v2

@@ Results.
We begin by assessing the need to mitigate noise in shallow circuits. The top panel of Fig. 2 shows the worst-case bias, $\max_{O,\varrho}\,\operatorname{bias}(O,\varrho)$, of standard non-mitigated shadow estimation for Pauli observables (see App. C for details) for different system sizes, circuit depths, and incoherent noise levels. The worst-case bias is monotonically increasing with all three parameters and the results are consistent with a linear increase of the bias with both the circuit depth and system size. For two-layer shadows, we observe Pauli worst-case biases that are already for three to four qubits about one order of magnitude larger than the average infidelity of the CNOTs used in the circuit. For example, targeting an estimation precision of $10^{-2}$, even fairly accurate CNOTs gates with infidelity $10^{-3}$ can introduce sizable biases in the linear estimation. Pauli observables with large non-trivial support tend to have larger biases, see App. D. We conclude that without mitigation using even very shallow shadows is typically not possible in practice already for small system sizes. This underpins the motivation to study our noise-robust variant.
@@
A key application of shadows with global Clifford rotations is fidelity estimation. Here, the sample complexity is constant in the system size, while randomized local basis measurements have exponential scaling [18]. This leads to the question of whether an ultra-shallow hardware-efficient ansatz can improve the sample efficiency even in the presence of experimental imperfections. In other words: Can the theoretically established sample-efficiency of classical shadows with entangling bases be translated into a practical advantage? We simulate the protocol of robust shadow fidelity estimation assuming perfectly prepared pure input states. The fidelity is estimated w.r.t. the input state itself. Only the two-qubit gates implementing the basis rotations for the shadow suffer from incoherent noise with average infidelity $10^{-3}$, and the number of samples is |$\absolutevalue{\mathcal{S}}=10^{5}$. The noisy frame operator is estimated with |$\absolutevalue{\mathcal{S}}=10^{6}$, which we found sufficient for ensuring that the estimation error of $\tilde{S}$ does not limit the accuracy of the fidelity estimation.
@@
Fig. 2 (bottom) shows the estimation error averaged over multiple uniformly random pure states for different system sizes and circuit depths. We find that the system-size dependence is significantly reduced already for shadows with $D=1,2$ when compared to local circuits. In effect, the estimation error is reduced by a factor of about $2$ already for these small systems of size $n=7,8$ when using short-depth circuits. Importantly, we observe that to harness these effects, the empirical estimation of $\tilde{S}$ put forward here is essential. Without noise mitigation, the noise-induced bias, depicted in the figure with dashed lines, deteriorates fidelity estimation by factors of $3$ to $4$ for the same small system sizes and $D=1$. For $D=2$ the estimation error using the noiseless frame operator $S$ is already almost one order of magnitude worse than the one that uses $\tilde{S}$.
Limitations of classical shadows in practice.
@@
In general, we expect that even assuming perfect bias mitigation, the improvements of the estimation error by going to deeper circuits will be limited by incoherent noise. Intuitively, every CNOT with incoherent errors will increase the probability that one only learns compromised information about the state. Even if the noise is perfectly mitigated by accurately estimating $\tilde{S}$, with some probability we learn less about the state. Thus, there will be an increased sample complexity of the fidelity estimation itself. At the same time, longer circuits (without noise) will improve the variance of fidelity estimation eventually reaching the efficiency of $3$-designs. We therefore have two competing effects on the sample complexity. We expect that for many estimation tasks, there exists a cross-over point depending on the infidelity of the entangling gates. Below this cross-over infidelity, increasing the depth improves estimation accuracy (at a fixed number of samples). Above it, accuracy is deteriorated by deeper circuits.
Figure 3: Tensor network representation of robust measurement frames of shallow circuits. (a) Worst-case bias for the estimation of Pauli observables introduced by converting $\tilde{S}$ to a matrix product state (TT) representation using sequential SVDs, relative to the size of the corresponding eigenvalue, as a function of $n$, for different bond dimensions $\chi$, and depths $D=1,2$. We observe that TT ranks $4$ and $8$ suffice to describe $\tilde{S}$ up to statistical error. (b) Worst-case bias for the estimation of Pauli observables when learning $\tilde{S}$ directly from the calibration data using a modified ALS method. We find a similar performance as for the sequential SVD. Colors indicate the permitted maximal ranks. Actual rank of the result of the rank-adaptive MALS is $\chi=4$ for $D=1$ ($\chi=8$ for $D=2$).
@@
We now take a more detailed look at the TT structure of the frame operator and its inverse. We begin with using sequential SVDs to solve the recovery problem in Eq. (6) for different ranks $\chi$ using the same noise setting as above. Fig. 3 panel (a) shows the additional (relative) worst-case bias in estimating Pauli observables when using a TT approximation to $f$ instead of the empirically learned dense vector. We indeed find that for $D=1$ a rank $\chi=4$ approximation yields a small relative error that is constant with the system size for $n\geq 6$ and higher ranks do not yield an improvement. Furthermore, we find that for smaller system sizes the MPS is also fitting the statistical fluctuations. For $D=2$ we find that $\chi=8$ accurately approximates $\tilde{S}$ up to statistical error.
We then apply the computationally and storage-efficient MALS algorithm to directly learn a TT representation of $\tilde{S}$ from the shadow data and calculate the induced bias in estimating Pauli observables. We initialize the algorithm with the analytical result for $S$ for the local Clifford frame. We observe an identical performance to the SVD. MALS is rank-adaptive but does not increase the rank beyond $\chi=8$ for $D=2$ even if permitted.
Figure 4: Simulation of overlap estimation from Ref. [5] under a gate-dependent noise model. (left) Worst-case bias for the estimation of Pauli observables introduced by implementing short-depth circuits with and without noise in an experiment, but using the ideal global Clifford frame in the post-processing stage, as a function of $D$, for two frames: $S_{n=4}^{\otimes 2}$ (noiseless, blue line) and $\tilde{S}_{n=4}^{\otimes 2}$ (noisy, brown line). Frames were estimated with |$\absolutevalue{\mathcal{S}}=10^{6}$. The noisy frame was generated under a gate-dependent circuit noise as well as readout noise. (right) Fidelity estimation error using aforementioned frames. States |$\ket{\Psi_{T}}$, |$\ket{\phi_{1}}$, and |$\ket{\phi_{2}}$ are defined in Appendix E.
[…]
@@
Quantum hardware demonstration.
In Fig. 5, we show the results of an experimental implementation of robust shallow shadows on the ibm_nairobi quantum processor from IBM Quantum. We implemented the measurement circuit of Fig. 1 in $5$ of the $7$ qubits available in the device, for depths $D\in\{0,1,2\}$. In Fig. 5 (a) we plot the estimated eigenvalues of $\tilde{S}$ as a function of the cardinality of the shadow $|\mathcal{S}|$ up to |$\absolutevalue{\mathcal{S}}=10^{5}$. (We also provide histograms for the distribution of samples in Fig. 4 (a) in App. F.) For simplicity, we show one eigenvalue per Pauli support. The dashed lines show the eigenvalues of the corresponding noiseless frame. In all three cases, the eigenvalues differ from their noiseless counterparts by a statistically significant margin. This confirms that the measurement frame to be inverted in practice should indeed be $\tilde{S}$ instead of $S$. We further observe that, as expected, with increasing depth the eigenvalues get closer. At $D=2$ the noiseless eigenvalues significantly deviate from each other and also the noise-induced deviation differs for different eigenvalues. Thus, an approximate construction with the (robust) global shadows is not permissible.
To cross-validate our results, we separate our $10^{5}$ samples in two batches: with a first batch of $90\%$ of the snapshots, we estimate $\tilde{S}$; with the remaining $10\%$, in turn, we perform a shadow tomography protocol on the $5$-qubit state |$\outerproduct{\mathbf{0}}{\mathbf{0}}$ | as well as on its marginal state |$\outerproduct{0}{0}$ | on one of the qubits. For each of the two states, we invert the classical shadows using both $\tilde{S}^{-1}$ and $S^{-1}$. In Fig. 5 (b), we plot the fidelity biases introduced by the mitigated and unmitigated estimations, respectively. For the unmitigated estimation, one can observe that at |$\absolutevalue{\mathcal{S}}\approx 10^{3}$ the statistical error becomes smaller than the bias (i.e. the systematic error due to the ‘wrong’ inverted frame). Increasing the number of samples does not improve the accuracy. In contrast, for the robust shallow shadows, the accuracy continues to increase steadily as |$\absolutevalue{\mathcal{S}}$ | goes beyond $10^{3}$.
Conclusions.
We presented a robust classical shadow scheme for a broad family of noisy, ultra-shallow measurement circuit ensembles. Under plausible noise assumptions, the measurement calibration stage admits an efficient tensor-network representation and can be performed using standard tensor-network tools. We illustrated the relevance of our method with both numerics and proof-of-principle experiments on an IBM Quantum device. Numerically, for relevant parameter regimes, we showed that while unmitigated shallow shadows with noisy circuits become more biased as the depth increases, robust ones become significantly more precise up to a crossing point. We thereby established that classical shadows can significantly improve sample efficiency in practice. At the same time, we established that the detrimental effect of incoherent noise on the sample efficiency fundamentally limits capitalizing on the efficiency of classical shadows in practice. Experimentally, we observed major error reductions in fidelity estimations. Our framework does not require analytical calculations of the effective measurement map but replaces it with an efficient direct estimation protocol. This offers versatile, practical estimation protocols that realize the efficacy of classical shadow as much as it is possible on actually existing and near-term quantum devices by using the available hardware-efficient noisy shallow circuits.
Note added.
During the completion of this paper, another independent work [36] appeared that also studies robust shallow shadows. After the publication of our preprint, Refs. [37] and [38] corroborated that global shadows can be de-randomized to log-depth. As we argue above, however, at current incoherent noise levels these constructions are not expected to improve the precision of shadow estimation and one still needs to resort to robust ultra-shallow shadows.
Acknowledgments.
We thank Stefano Carrazza and Stavros Efthymiou for their insights on numerical simulations. We further thank Michał Oszmaniec for discussions on noise mitigation. We thank Jadwiga Wilkens for the availability of the Quantum Circuit Library [39].
Appendix A Details on tensor network description
$\begin{split}\operatorname{CNOT}&=\sum_{j\in\{0,1\}}\outerproduct{j}{j}\otimes X^{j}\\ &=\left(\sum_{j\in\{0,1\}}\outerproduct{j}{j}\otimes\bra{j}\otimes\mathbbm{1}\right)\left(\mathbbm{1}\otimes\sum_{l\in\{0,1\}}\ket{l}\otimes X^{l}\right)\\ &=(L\otimes\mathbbm{1})(\mathbbm{1}\otimes R)\,,\end{split}$ (7)
where $X$ denotes the Pauli-$X$ gate, and the last line defines the operators $L$ and $R$. Now we define the MPO core $C_{j,l}\coloneqq(\bra{j}\otimes\mathbbm{1})R\,L(\mathbbm{1}\otimes\ket{l})\in\mathbb{C}^{2\times 2}$. The unitary operator $L_{\operatorname{CNOT}}$ of a $1$D layer of CNOT gates on $n$ qubits with periodic boundary conditions and even $n$ can be written as
@@ Appendix A Details on tensor network description
The operator $\tilde{S}$ can also be written as a shallow circuit of operators acting on quantum channels. Assuming local invariance, also $f=(f_{k})_{k\in[n]}\in\mathbb{R}^{n}$ has a corresponding circuit representation. The rank of the resulting TT description of $f$ is increased by every layer of entangling gates and non-local noise. For fairly general noise affecting the shallow circuits, $f$ is of low TT rank. To this end, it is instructive to discuss a scenario with strictly local and mostly gate-dependent noise. Let us assume that $\Omega^{(l)}_{j}(U)$ is the CPT map describing the noisy implementation of the gate $U\in\mathrm{U}(2)$ when applied on qubit $l$ in the circuit layer $j$. This noise model is general enough to also capture local noise and errors affecting the read-out and entangling gates. These noise effects can be included in the respective neighboring layer of single-qubit rotations. The local channel twirl with a group $G$ including noise is formally described by the non-commutative Fourier transform of the map $\Omega^{(l)}_{j}$ evaluated at the representation of unitary channels $\omega(U)=U\otimes\bar{U}$, i.e.
$\hat{\Omega}^{(l)}_{j}[\omega]=\int_{G}\bar{U}\otimes U\otimes\Omega^{(l)}_{j}(U)\mathrm{d}\mu(U)\,.$ (9)
We will here simply use $\hat{\Omega}^{(l)}_{j}[\omega]$ as a convenient symbol for the “noisy channel twirl”. Note that $\hat{\Omega}^{(l)}_{j}[\omega]$ is a linear operator on superoperators, a “super-duper-operator” if one insists. We can also introduce a corresponding operator for the entangling layer that conjugates a quantum channel with unitary channels of a layer of CNOT gates. The resulting MPO, $\mathcal{C}=\bar{L}_{\operatorname{CNOT}}\otimes L_{\operatorname{CNOT}}\otimes L_{\operatorname{CNOT}}\otimes\bar{L}_{\operatorname{CNOT}}$, has bond dimension $16$. The read-out is locally described by $\mathcal{B}_{2}=\sum_{z\in[2]}{|\Pi_{z})}{(\Pi_{z}|}$, the local dephasing channel in computational basis. Finally, to ensure the local invariance of $\tilde{S}$, we assume gate-independent left noise $\Omega^{(1)}_{j}(U)=\Lambda_{j}\omega(U)$ for the first layer acting on the state under consideration, with $\Lambda_{j}$ being a quantum channel.
Figure 7 depicts the resulting expression for $f$. Without any further assumption, we conclude that $f$ has a TT rank of at most $16^{D}$. The rank is even smaller if all noise channels are Pauli noise. Then, we can fully restrict the circuit in Fig. 7 as acting only on the subspace of Pauli noise channels, i.e.  diagonal channels in Pauli operator basis. To see this note that the local dephasing channel is a Pauli noise channel. Under the assumption of Pauli noise, the noisy channel twirl maps Pauli noise channels to Pauli noise channels. Being a Clifford unitary, the entangling layer acts as a two-local permutation on the diagonal of the Pauli-noise channel. The representation of $\mathcal{C}$ restricted to Pauli-noise channels has (MPO) bond dimension $4$. Hence, $f$ is described by a TT of rank at most $\chi=4^{D}$.
The argument for one-local noise can be generalized to quasi-local noise by introducing, e.g. , MPO descriptions of the implementation map modeling cross-talk effects, or modeling a noisy entangling layer $\mathcal{C}$ with a potentially larger bond dimension. In general, the TT rank of $f$ will be increased due to correlated noise effects. If these correlations are sufficiently short-ranged, $f$ still has an effective low-rank description. We expect that this will often be the case on actual quantum hardware.
Figure 7: Tensor network of $f$ for strictly local noise mostly gate-dependent noise. The entangling layer acting here as a “super-duper operator” on quantum channels has an MPO bond dimension $16$, and restricted to Pauli noise channels it has an MPO bond dimension $4$. Correspondingly, the TT rank of $f$ is $16^{D}$ for local noise and $4^{D}$ for local Pauli noise.
Appendix B Incoherent noise model
@@
Appendix C Worst-case bias measure
@@
where we used the diagonality of $\tilde{S}$ in Pauli basis and that $|(w_{a}\,|\,\varrho)|\leq\|w_{a}\|_{\infty}\|\varrho\|_{1}\leq 1/\sqrt{d}$ for all $a,\varrho$, and $\|\cdot\|_{1}$ denoting the trace-norm (nuclear / Schatten-$1$ norm). Most importantly, this bound is saturated for $\varrho$ in an eigenstate of $w_{a}$ the observable suffering the maximal bias.